# Dual metrics on the boundary of strictly polyhedral hyperbolic 3-manifolds

### v4

@@ 3 Infinitesimal rigidity
Let $\gamma$ be a smooth simple closed curve on $\Xi$. Orient it arbitrarily and pick a smooth collar $\bar{C}_{\gamma}\subset\Xi$ to the left of $\gamma$ parametrized as $\gamma\times[0,1]$. Let $\phi:[0,1]\rightarrow[0,1]$ be a smooth function equal 0 in a neighborhood of zero and 1 in a neighborhood of one. Let $\phi_{\gamma}:\bar{C}_{\gamma}\rightarrow\mathbb{R}_{+}$ be a smooth function that is equal to $\phi$ on every $[0,1]$-fiber of the parametrization of $\bar{C}_{\gamma}$. For $p\in\gamma$ we pick an identification of the fiber of $E$ over $p$ with $\mathfrak{g}$ and choose there the element $F(\rho(\gamma))$. Since $F(\rho(\gamma))$ is ${\rm Ad}(\rho(\gamma))$-invariant, it defines a parallel section $\sigma_{\gamma}$ over $\bar{C}_{\gamma}$. Define $\omega_{\gamma}:=d\phi_{\gamma}\otimes\sigma_{\gamma}\in\Omega^{1}_{L^{2}}(\Xi,E)$, which is a smooth compactly supported 1-form. Since $\sigma_{\gamma}$ is a parallel section, $\omega_{\gamma}$ is closed. Denote its class by $y_{\gamma}\in H^{1}_{L^{2}}(\Xi,E)$. A simple computation using the Stokes Theorem shows that
@@
for every $y\in H^{1}_{L^{2}}(\Xi,E)$, see [45, p. 378].
We pick disjoint collars $\bar{C}_{\gamma}$ for all $\gamma\in\mathcal{P}$ and produce the 1-forms $\omega_{\gamma}$ by the above procedure. Since $\{d{\rm tr}_{\gamma}\}_{\gamma\in\mathcal{P}}$ are $\mathbb{C}$-linearly independent in $H^{1}_{L^{2}}(\Xi,E)^{*}$, (3.5) implies that the classes $y_{\gamma}$ are $\mathbb{C}$-linearly independent in $H^{1}_{L^{2}}(\Xi,E)$. Further, the disjointness of the supports of $\omega_{\gamma}$, the skew-symmetry of $H$, (3.4) and (3.5) mean that $y_{\gamma}\in\bigcap_{\gamma\in\mathcal{P}}\ker(d{\rm tr}_{\gamma})$ for every $\gamma\in\mathcal{P}$. This and (3.3) result in
@@
Furthermore, [71, Corollary 2.3 and Lemma 3.14] imply the $L^{2}$-Hodge Theorem
@@
where $\varepsilon(.)$ and $\iota(.)$ are the extrinsic and intrinsic products on forms. Then $d^{E}=D+T$ and $\delta^{E}=D^{t}+T^{t}$. Note that $T$ and $T^{t}$ are bounded 0th-order operators, see details in [30, 39, 70, 71].
@@
$\|\xi(r)\|_{L^{2}(\Xi_{r})}=O(r^{\alpha+1}),~{}~{}~{}~{}~{}\|(\nabla_{e_{i}}\xi)(r)\|_{L^{2}(\Xi_{r})}=O(r^{\alpha})~{}~{}~{}~{}~{}{\rm for}~{}i=1,2.$
@@
Since $y\in\bigcap_{e}\ker(d{\rm tr}_{\gamma_{e}})$, Lemma 3.7 implies that $y\in H_{L^{2}}(L,E)$. By Lemma 3.8 it can be represented by an $L^{2}$-harmonic 1-form $\omega\in\Omega^{1}_{L^{2}}(L,E)$. Define $L_{r}:=L\backslash B_{r}(\Sigma)$ and $K_{r}:=\partial L_{r}$. Integrating by parts, we obtain (see [30, Proposition 1.3])
@@
We decompose $K_{r}$ into the union of $K_{e,r}:=\partial B_{r}(e)\backslash(\bar{B}_{r}(u_{1})\cup\bar{B}_{r}(u_{2}))$ where $e$ is an edge of $\Sigma$ with vertices $u_{1}$ and $u_{2}$, and of $K_{u,r}:=K_{r}\cap\partial B_{r}(u)$ where $u$ is a vertex of $\Sigma$. We denote the respective integrals over $K_{e,r}$ and $K_{u,r}$ by $A_{e,r}$ and $A_{u,r}$. The terms $A_{e,r}$ are bounded exactly as in [71, Proposition 3.17] without changes. What is different is the bound on $A_{u,r}$.
At every point $p\in L$ the fiber $E_{p}$ can be decomposed into the direct orthogonal sum of the infinitesimal rotations around $p$ and the infinitesimal translations. The second summand is naturally isomorphic to $T_{p}L$. A special feature of dimension 3 is that the first summand is also naturally isomorphic to $T_{p}L$, see [30, Section 2]. We denote the respective decomposition of $E$ by $E=E^{1}\oplus E^{2}$. Decompose respectively $\omega=(\omega^{1},\omega^{2})$. Then
@@
Substituting this into (3.6), we decompose $A_{u,r}$ into two terms, $A_{u,r}=A_{u,r}^{s}+A_{u,r}^{\omega}$, where $A_{u,r}^{\omega}:=2\int_{K_{u,r}}\omega^{1}_{u}\wedge\omega^{2}_{u}$. Let us bound the first summand. Let $s=(s^{1},s^{2})$ according to $E=E^{1}\oplus E^{2}$. Because $\omega$ is coclosed, we get $\Delta^{E}s=-\delta^{E}\omega_{u}$, where $\Delta^{E}$ is the Hodge Laplacian of the bundle $E$. We claim that $\delta^{E}\omega_{u}$ is in $L^{2}$. Indeed, the holonomy restricted to $\Xi_{u}$ preserves a point $p\in\mathbb{H}^{3}$. The bundle $E$ over $B_{r_{0}}(u)$ can be endowed with a parallel metric inherited from $T_{p_{0}}\mathbb{H}^{3}$ (this is not the standard metric of $E$). Denote by $\delta^{E}_{0}$ the codifferential on $\Omega^{\bullet}(L,E)$ with respect to the new metric. Note that on $\Xi_{u}$ the section $\delta^{E}_{0}\omega_{u}$ is compactly supported (because $\omega_{u}$ on $\Xi_{u}$ is compactly supported), thus it is in $L^{2}$. It follows that $\delta^{E}_{0}\omega_{u}$ on $B_{r_{0}}(u)$ is also in $L^{2}$. But the difference $\delta^{E}-\delta^{E}_{0}$ is a bounded operator, see [71, p. 353]. Since $\omega_{u}$ is itself in $L^{2}$ on $B_{r_{0}}(u)$, we get that $\delta^{E}\omega_{u}$ is in $L^{2}$ as claimed.
As we mentioned, both sub-bundles $E^{1}$ and $E^{2}$ are naturally isomorphic to $TL$. Hence, we may view $s^{1}$ and $s^{2}$ as vector fields on $L$. Denote by $\xi^{1}$ and $\xi^{2}$ the real-valued 1-forms dual to to $s^{1}$ and $s^{2}$ and by $\zeta^{1}$ and $\zeta^{2}$ the real-valued 1-forms dual to the components of $-\delta^{E}\omega_{u}$. The equation $\Delta^{E}s=-\delta^{E}\omega_{u}$ translates to (see [71, p. 347])
@@
$\|\xi^{j}(r)\|_{L^{2}(\Xi_{u,r})}=O(r^{\alpha+1}),~{}~{}~{}~{}~{}\|(\nabla_{e_{i}}\xi^{j})(r)\|_{L^{2}(\Xi_{u,r})}=O(r^{\alpha}),$ (3.7)
@@
We expand this expression into further integrals of the scalar products of the individual summands and bound every integral separately by the Cauchy–Schwartz Inequality. We will compute the $L^{2}$-norms not over $K_{u,r}$, but rather over $\Xi_{u,r}\supset K_{u,r}$. For terms $\omega^{j}_{u}(e_{i})$ we use the trivial pointwise bounds $|\omega^{j}_{u}(e_{i})|=O(r^{-1})$, thus $\|\omega^{j}_{u}(e_{i})\|_{L^{2}(\Xi_{u,r})}=O(1)$ as $r\rightarrow 0$. Together with (3.7) and with the fact that $T$ is bounded it gives $|A_{u,r}^{s}|=O(r^{\alpha})$.
@@
For a given $\gamma\in\mathcal{P}_{u}$ the form $\omega_{\gamma}$ on $\Xi_{u}$ is supported in $\bar{C}_{\gamma}$, which belongs to one of the pair of pants determined by $\mathcal{P}_{u}$ (since the collars are disjoint for different curves). Let us denote this pair of pants by $\Xi^{\prime}_{\gamma}$. Denote the respective surface $\Xi(\alpha_{1},\alpha_{2},\alpha_{3})$, given by the suitability of $\mathcal{P}_{u}$, by $\Xi_{\gamma}$. Hence, after a scaling, $\Xi^{\prime}_{\gamma}$ naturally embeds to $\Xi_{\gamma}$. The spherical metric provides us a holonomy map $\rho_{\gamma}:\pi_{1}(\Xi_{\gamma})\rightarrow SO(2)$. We consider $SO(2)$ as a subgroup of $SL(2,\mathbb{C})=\widetilde{G}$, and view $\rho_{\gamma}$ as having values in $\widetilde{G}$. Thus, we can define the twisted ${\rm Ad}\circ\rho_{\gamma}$-bundle over $\Xi_{\gamma}$ with the fiber $\mathfrak{g}$, which we denote by $E_{\gamma}$. Since such a bundle depends only on the conjugacy class of a representation, the pullback of $E_{\gamma}$ to $\Xi^{\prime}_{\gamma}$ is naturally isomorphic to the restriction of $E$ to $\Xi^{\prime}_{\gamma}$. We now denote $E_{\gamma}$ also just by $E$ and consider the 1-form $\omega_{\gamma}$ now as a $E$-valued 1-form on $\Xi_{\gamma}$. We extend $E$ to the bundle over $B_{\gamma,r_{0}}:=(0,r_{0})\times\Xi_{\gamma}$ equipped with the hyperbolic cone-metric. The radial projection of $B_{\gamma,r_{0}}$ to $\Xi_{\gamma}$ is a homotopy equivalence, and we continue to denote the pullback of $\omega_{\gamma}$ to $B_{\gamma,r_{0}}$ by $\omega_{\gamma}$. Define $C_{\gamma,r}$ to be the intersection of the support of $\omega_{\gamma}$ with $\{r\}\times\Xi_{\gamma}$.
Since $\omega_{\gamma}$ is compactly supported on $\Xi_{\gamma}$, it is in $L^{2}$, both on $\Xi_{\gamma}$ and on $B_{\gamma,r_{0}}$. Due to [71, Lemma 3.9], we have $H^{1}_{L^{2}}(\Xi_{\gamma},E)=0$. Thereby, we have $\omega_{\gamma}=d^{E}s$ for a smooth $L^{2}$-section $s$ on $\Xi_{\gamma}$, which we extend to $B_{\gamma,r_{0}}$. We get $\Delta^{E}s=-\delta^{E}\omega_{\gamma}$ on $B_{\gamma,r_{0}}$. Pick an orthonormal frame $(e_{1},e_{2},e_{3})$, $e_{3}=\partial/\partial r$, on $B_{\gamma,r_{0}}$. We have
@@ 4 Properness
endowed with the metric induced from .,.$\langle.,.\rangle_{E}$. The central projection from the origin of $\mathbb{R}^{4}$ maps injectively $\mathbb{H}^{3}$ and $d\mathbb{S}^{3}$ to subsets of $\mathbb{S}^{3}$ denoted by $\mathbb{H}_{s}^{3}$ and $d\mathbb{S}_{s}^{3}$. Notably it sends geodesics to geodesics, and, thus, convex sets to convex sets. If $C\subset\mathbb{H}_{s}^{3}$ or $C\subset d\mathbb{S}_{s}^{3}$ is (future-)convex, then the dual set $C^{*}$ can be obtained from the usual polarity on $\mathbb{S}^{3}$ composed with the reflection with respect to the hyperplane $x_{0}=0$. The boundary at infinity $\partial_{\infty}d\mathbb{S}^{3}$ projects to two spherical caps denoted by $\partial_{\infty}^{+}d\mathbb{S}_{s}^{3}$ and $\partial_{\infty}^{-}d\mathbb{S}_{s}^{3}$.
@@
But this is not the case for us, because the metric $d$ is complete. Indeed, we claim that all limit points of $\phi(D)$ that do not belong to $\phi(D)$ are in $\partial_{\infty}^{+}d\mathbb{S}^{3}$. Suppose the contrary, that such a point $q$ exists in $d\mathbb{S}^{3}$. Then $q\in\partial{\rm conv}_{s}(\phi(D))$. Pick $p\in\phi(D)$, there exists a rectifiable curve $\gamma:[0,a_{0}]\rightarrow\partial{\rm conv}_{s}(D)\cap d\mathbb{S}^{3}$ connecting $p$ and $q$. Let $[0,a]$ be the maximal segment such that $\gamma|_{[0,a)}$ belongs to $\phi(D)$. Since $\phi(D)$ is relatively open in $\partial{\rm conv}_{s}(\phi(D))$, the point $\gamma(a)$ is not in $\phi(D)$. But $\phi^{-1}\circ\gamma|_{[0,a)}$ is a curve of finite length in $(D,d)$, hence, due to completeness of $d$, it extends to $a$. Thereby, $\gamma(a)\in\phi(D)$, which is a contradiction. It follows that ${\rm conv}_{s}(\phi(D))\cap d\mathbb{S}^{3}_{s}$ is a future-convex set and
@@
We denote this point $\phi(v)$, so $\phi$ is a map from $\widetilde{W}$ to $\overline{d}\mathbb{S}_{s}^{3}$.
@@
Define $X:=D\backslash{\rm ost}(\phi^{-1}(\partial_{\infty}^{+}d\mathbb{S}_{s}^{3}))$. By construction, $X$ is non-empty. By continuity, the restriction of $\phi$ to $X$ is a path-isometry. We claim that the restriction of $\phi$ to $X$ is injective. Indeed, since it is a path-isometry, it is locally injective. Now, suppose that for two points $x,y\in X$ we have $\phi(x)=\phi(y)$. Due to local injectivity, $x\notin{\rm ost}(y)$ and vice versa. Pick some small $r$ such that for all sufficiently large $i$ the $r$-neighborhood of $y$ in $(D,d_{i})$ is contained in ${\rm ost}(y)$. As $\phi(x)\in d\mathbb{S}^{3}_{s}$, we can pick its neighborhood $U$ such that (1) for every point $p\in U$ in space-like relation to $\phi(x)$ the future-cones of $\phi(x)$ and of $p$ intersect and so do their past-cones, and (2) the length of any space-like segment in $U$ is at most $r/2$. Next, for every $i$ we pick a neighborhood $U_{i}$ of $\phi_{i}(x)$ converging to $U$ satisfying the same conditions, except that in (2) we replace $r/2$ with $r$. As $\phi(x)=\phi(y)$, for large enough $i$ we get $\phi_{i}(y)\in U_{i}$. For $p\in d\mathbb{S}^{3}$ by $FP(p)$ denote the union of the closures of its future- and past-cone. Then $\phi_{i}({\rm ost}(y))$ intersects $FP(\phi_{i}(x))$. Indeed, by (1) the space-like cone from $\phi_{i}(y)$ spanned by $\phi_{i}({\rm ost}(y))$ intersects $FP(\phi_{i}(x))$, and by (2) there is an intersection point belonging to $\phi_{i}({\rm ost}(y))$. This is a contradiction with $\phi_{i}$ being future-convex.

### v5

@@ 3 Infinitesimal rigidity
Let $\gamma$ be a smooth simple closed curve on $\Xi$. Orient it arbitrarily and pick a smooth collar $\bar{C}_{\gamma}\subset\Xi$ to the left of $\gamma$ parametrized as $\gamma\times[0,1]$. Let $\phi:[0,1]\rightarrow[0,1]$ be a smooth function equal to 0 in a neighborhood of zero and to 1 in a neighborhood of one. Let $\phi_{\gamma}:\bar{C}_{\gamma}\rightarrow\mathbb{R}_{+}$ be a smooth function that is equal to $\phi$ on every $[0,1]$-fiber of the parametrization of $\bar{C}_{\gamma}$. For $p\in\gamma$ we pick an identification of the fiber of $E$ over $p$ with $\mathfrak{g}$ and choose there the element $F(\rho(\gamma))$. Since $F(\rho(\gamma))$ is ${\rm Ad}(\rho(\gamma))$-invariant, it defines a parallel section $\sigma_{\gamma}$ over $\bar{C}_{\gamma}$. Define $\omega_{\gamma}:=d\phi_{\gamma}\otimes\sigma_{\gamma}\in\Omega^{1}_{L^{2}}(\Xi,E)$, which is a smooth compactly supported 1-form. Since $\sigma_{\gamma}$ is a parallel section, $\omega_{\gamma}$ is closed. Denote its class by $y_{\gamma}\in H^{1}_{L^{2}}(\Xi,E)$. A simple computation using the Stokes Theorem shows that
@@
for every $y\in H^{1}_{L^{2}}(\Xi,E)$, see [47, p. 378].
We pick disjoint collars $\bar{C}_{\gamma}$ for all $\gamma\in\mathcal{P}$ and produce the 1-forms $\omega_{\gamma}$ by the above procedure. Since $\{d{\rm tr}_{\gamma}\}_{\gamma\in\mathcal{P}}$ are $\mathbb{C}$-linearly independent in $H^{1}_{L^{2}}(\Xi,E)^{*}$, (3.5) implies that the classes $y_{\gamma}$ are $\mathbb{C}$-linearly independent in $H^{1}_{L^{2}}(\Xi,E)$. Further, the disjointness of the supports of $\omega_{\gamma}$, the skew-symmetry of $H$, (3.4) and (3.5) mean that $y_{\gamma}\in\bigcap_{\gamma^{\prime}\in\mathcal{P}}\ker(d{\rm tr}_{\gamma^{\prime}})$ for every $\gamma\in\mathcal{P}$. This and (3.3) result in
@@
Furthermore, [74, Corollary 2.3 and Lemma 3.14] imply the $L^{2}$-Hodge Theorem
@@
where $\varepsilon(.)$ and $\iota(.)$ are the extrinsic and intrinsic products on forms, and $E_{i}$ is the infinitesimal translation in the direction $e_{i}$. Then $d^{E}=D+T$ and $\delta^{E}=D^{t}+T^{t}$. Note that $T$ and $T^{t}$ are bounded 0th-order operators, see details in [32, 41, 73, 74].
@@
$\|\xi\|_{L^{2}(\Xi_{r})}=O(r^{\alpha+1}),~{}~{}~{}~{}~{}\|\nabla_{e_{i}}\xi\|_{L^{2}(\Xi_{r})}=O(r^{\alpha})~{}~{}~{}~{}~{}{\rm for}~{}i=1,2.$
Here the norms are taken for the restrictions of the respective forms to $\Xi_{r}$.
@@
Since $y\in\bigcap_{e}\ker(d{\rm tr}_{\gamma_{e}})$, Lemma 3.7 implies that $y\in H_{L^{2}}(L,E)$. By Lemma 3.8 it can be represented by an $L^{2}$-harmonic 1-form $\omega\in\Omega^{1}_{L^{2}}(L,E)$. Define $L_{r}:=L\backslash B_{r}(\Sigma)$ and $K_{r}:=\partial L_{r}$. Integrating by parts, we obtain (see [32, Proposition 1.3])
@@
We decompose $K_{r}$ into the union of $K_{e,r}:=\partial B_{r}(e)\backslash(\bar{B}_{r}(u_{1})\cup\bar{B}_{r}(u_{2}))$ where $e$ is an edge of $\Sigma$ with vertices $u_{1}$ and $u_{2}$, and of $K_{u,r}:=K_{r}\cap\partial B_{r}(u)$ where $u$ is a vertex of $\Sigma$. We denote the respective integrals over $K_{e,r}$ and $K_{u,r}$ by $A_{e,r}$ and $A_{u,r}$. The terms $A_{e,r}$ are bounded exactly as in [74, Proposition 3.17] without changes. What is different is the bound on $A_{u,r}$.
At every point $p\in L$ the fiber $E_{p}$ can be decomposed into the direct orthogonal sum of the infinitesimal rotations around $p$ and the infinitesimal translations. The second summand is naturally isomorphic to $T_{p}L$. A special feature of dimension 3 is that the first summand is also naturally isomorphic to $T_{p}L$, see [32, Section 2]. We denote the respective decomposition of $E$ by $E=E^{1}\oplus E^{2}$. Decompose respectively $\omega=(\omega^{1},\omega^{2})$. Then
@@
Substituting this into (3.6), we decompose $A_{u,r}$ into two terms, $A_{u,r}=A_{u,r}^{s}+A_{u,r}^{\omega}$, where $A_{u,r}^{\omega}:=2\int_{K_{u,r}}\omega^{1}_{u}\wedge\omega^{2}_{u}$. Let us bound the first summand. Let $s=(s^{1},s^{2})$ according to $E=E^{1}\oplus E^{2}$. Because $\omega$ is coclosed, we get $\Delta^{E}s=-\delta^{E}\omega_{u}$, where $\Delta^{E}$ is the Hodge Laplacian of the bundle $E$. We claim that $\delta^{E}\omega_{u}$ is in $L^{2}$. Indeed, the holonomy restricted to $\Xi_{u}$ preserves a point $p\in\mathbb{H}^{3}$. The bundle $E$ over $B_{r_{0}}(u)$ can be endowed with a parallel metric inherited from $T_{p}\mathbb{H}^{3}$ (this is not the standard metric of $E$). Denote by $\delta^{E}_{0}$ the codifferential on $\Omega^{\bullet}(B_{r_{0}}(u),E)$ with respect to the new metric. Note that on $\Xi_{u}$ the section $\delta^{E}_{0}\omega_{u}$ is compactly supported (because $\omega_{u}$ on $\Xi_{u}$ is compactly supported), thus it is in $L^{2}$. It follows that $\delta^{E}_{0}\omega_{u}$ on $B_{r_{0}}(u)$ is also in $L^{2}$. But the difference $\delta^{E}-\delta^{E}_{0}$ is a bounded operator, see [74, p. 353]. Since $\omega_{u}$ is itself in $L^{2}$ on $B_{r_{0}}(u)$, we get that $\delta^{E}\omega_{u}$ is in $L^{2}$ as claimed.
As we mentioned, both sub-bundles $E^{1}$ and $E^{2}$ are naturally isomorphic to $TL$. Hence, we may view $s^{1}$ and $s^{2}$ as vector fields on $L$. Denote by $\xi^{1}$ and $\xi^{2}$ the real-valued 1-forms dual to to $s^{1}$ and $s^{2}$ and by $\zeta^{1}$ and $\zeta^{2}$ the real-valued 1-forms dual to the components of $-\delta^{E}\omega_{u}$. The equation $\Delta^{E}s=-\delta^{E}\omega_{u}$ translates to (see [74, p. 347])
@@
$\|\xi^{j}\|_{L^{2}(\Xi_{u,r})}=O(r^{\alpha+1}),~{}~{}~{}~{}~{}\|\nabla_{e_{i}}\xi^{j}\|_{L^{2}(\Xi_{u,r})}=O(r^{\alpha}),$ (3.7)
@@
We expand this expression into further integrals of the scalar products of the individual summands and bound every integral separately by the Cauchy–Schwarz Inequality. We will compute the $L^{2}$-norms not over $K_{u,r}$, but rather over $\Xi_{u,r}\supset K_{u,r}$. For terms $\omega^{j}_{u}(e_{i})$ we use the trivial pointwise bounds $|\omega^{j}_{u}(e_{i})|=O(r^{-1})$, thus $\|\omega^{j}_{u}(e_{i})\|_{L^{2}(\Xi_{u,r})}=O(1)$ as $r\rightarrow 0$. Together with (3.7) and with the fact that $T$ is bounded it gives $|A_{u,r}^{s}|=O(r^{\alpha})$.
@@
For a given $\gamma\in\mathcal{P}_{u}$ the form $\omega_{\gamma}$ on $\Xi_{u}$ is supported in $\bar{C}_{\gamma}$, which belongs to one of the pair of pants determined by $\mathcal{P}_{u}$ (since the collars are disjoint for different curves). Let us denote this pair of pants by $\Xi^{\prime}_{\gamma}$. Denote the respective surface $\Xi(\alpha_{1},\alpha_{2},\alpha_{3})$, given by the suitability of $\mathcal{P}_{u}$, by $\Xi_{\gamma}$. Hence, after a scaling, $\Xi^{\prime}_{\gamma}$ naturally embeds to $\Xi_{\gamma}$. The spherical metric provides us a holonomy map $\rho_{\gamma}:\pi_{1}(\Xi_{\gamma})\rightarrow{\rm SO}(2)$. We consider ${\rm SO}(2)$ as a subgroup of ${\rm SL}(2,\mathbb{C})=\widetilde{G}$, and view $\rho_{\gamma}$ as having values in $\widetilde{G}$. Thus, we can define the twisted ${\rm Ad}\circ\rho_{\gamma}$-bundle over $\Xi_{\gamma}$ with the fiber $\mathfrak{g}$, which we denote by $E_{\gamma}$. Since such a bundle depends only on the conjugacy class of a representation, the pullback of $E_{\gamma}$ to $\Xi^{\prime}_{\gamma}$ is naturally isomorphic to the restriction of $E$ to $\Xi^{\prime}_{\gamma}$. We now denote $E_{\gamma}$ also just by $E$ and consider the 1-form $\omega_{\gamma}$ now as a $E$-valued 1-form on $\Xi_{\gamma}$. We extend $E$ to the bundle over $B_{\gamma,r_{0}}:=(0,r_{0})\times\Xi_{\gamma}$ equipped with the hyperbolic cone-metric. The radial projection of $B_{\gamma,r_{0}}$ to $\Xi_{\gamma}$ is a homotopy equivalence, and we continue to denote the pullback of $\omega_{\gamma}$ to $B_{\gamma,r_{0}}$ by $\omega_{\gamma}$. Define $C_{\gamma,r}$ to be the intersection of the support of $\omega_{\gamma}$ with $\{r\}\times\Xi_{\gamma}$.
Since $\omega_{\gamma}$ is compactly supported on $\Xi_{\gamma}$, it is in $L^{2}$, both on $\Xi_{\gamma}$ and on $B_{\gamma,r_{0}}$. Due to [74, Lemma 3.9], we have $H^{1}_{L^{2}}(\Xi_{\gamma},E)=0$. Thereby, we have $\omega_{\gamma}=d^{E}s$ for a smooth $L^{2}$-section $s$ on $\Xi_{\gamma}$, which we extend to $B_{\gamma,r_{0}}$. We get $\Delta^{E}s=\delta^{E}\omega_{\gamma}$ on $B_{\gamma,r_{0}}$. Pick an orthonormal frame $(e_{1},e_{2},e_{3})$, $e_{3}=\partial/\partial r$, on $B_{\gamma,r_{0}}$. We have
@@ 4 Properness
endowed with the metric induced from .,.$\langle.,.\rangle_{E}$. The central projection from the origin of $\mathbb{R}^{4}$ maps injectively $\mathbb{H}^{3}$ and $d\mathbb{S}^{3}$ to subsets of $\mathbb{S}^{3}$ denoted by $\mathbb{H}_{s}^{3}$ and $d\mathbb{S}_{s}^{3}$. Notably it sends geodesics to geodesics, and, thus, convex sets to convex sets. The boundary at infinity $\partial_{\infty}d\mathbb{S}^{3}$ projects to two spherical caps denoted by $\partial_{\infty}^{+}d\mathbb{S}_{s}^{3}$ and $\partial_{\infty}^{-}d\mathbb{S}_{s}^{3}$.
Remark 4.2.
If $C\subset\mathbb{H}_{s}^{3}$ or $C\subset d\mathbb{S}_{s}^{3}$ is (future-)convex, then the dual set $C^{*}$ can be obtained from the usual polarity on $\mathbb{S}^{3}$ composed with the reflection with respect to the hyperplane $x_{0}=0$. Indeed, this follows from the fact that the linear plane, orthogonal to a vector with respect to the Minkowski metric, differs from the linear plane, orthogonal to the same vector with respect to the Euclidean metric, by reflection with respect to the hyperplane $x_{0}=0$.
@@
But this is not the case for us, because the metric $d$ is complete. Indeed, we claim that all limit points of $\phi(D)$ that do not belong to $\phi(D)$ are in $\partial_{\infty}^{+}d\mathbb{S}^{3}$. Suppose the contrary, that such a point $q$ exists in $d\mathbb{S}^{3}$. Then $q\in\partial{\rm conv}_{s}(\phi(D))$. Pick $p\in\phi(D)$, there exists a rectifiable curve (with respect to the spherical metric) $\gamma:[0,a_{0}]\rightarrow\partial{\rm conv}_{s}(D)\cap d\mathbb{S}^{3}$ connecting $p$ and $q$. Let $[0,a]$ be the maximal segment such that $\gamma|_{[0,a)}$ belongs to $\phi(D)$. The curve $\gamma|_{[0,a)}$ has finite spherical length and belongs to a space-like polyhedral surface, hence it also has a finite de Sitter length. Since $\phi(D)$ is relatively open in $\partial{\rm conv}_{s}(\phi(D))$, the point $\gamma(a)$ is not in $\phi(D)$. But $\phi^{-1}\circ\gamma|_{[0,a)}$ is a curve of finite length in $(D,d)$, hence, due to completeness of $d$, it extends to $a$. Thereby, $\gamma(a)\in\phi(D)$, which is a contradiction. It follows that ${\rm conv}_{s}(\phi(D))\cap d\mathbb{S}^{3}_{s}$ is a future-convex set and
@@
We denote this point by $\phi(v)$, so $\phi$ is a map from $\widetilde{W}$ to $\overline{d}\mathbb{S}_{s}^{3}$.
@@
Define $X:=D\backslash{\rm ost}(\phi^{-1}(\partial_{\infty}^{+}d\mathbb{S}_{s}^{3}))$. By construction, $X$ is non-empty. By continuity, the restriction of $\phi$ to $X$ is a path-isometry. We claim that the restriction of $\phi$ to $X$ is injective. Indeed, since it is a path-isometry, it is locally injective. Now suppose that for two points $x,y\in X$ we have $\phi(x)=\phi(y)$. Due to local injectivity, $x\notin{\rm ost}(y)$ and vice versa. Pick some small $r$ such that for all sufficiently large $i$ the $r$-neighborhood of $y$ in $(D,d_{i})$ is contained in ${\rm ost}(y)$. As $\phi(x)\in d\mathbb{S}^{3}_{s}$, we can pick its neighborhood $U$ such that (1) for every point $p\in U$ in space-like relation to $\phi(x)$ the future-cones of $\phi(x)$ and of $p$ intersect and so do their past-cones, and (2) the length of any space-like segment in $U$ is at most $r/2$. Next, for every $i$ we pick a neighborhood $U_{i}$ of $\phi_{i}(x)$ converging to $U$, satisfying the same conditions, except that in (2) we replace $r/2$ with $r$. As $\phi(x)=\phi(y)$, for large enough $i$ we get $\phi_{i}(y)\in U_{i}$. For $p\in d\mathbb{S}^{3}$ by $FP(p)$ denote the union of the closures of its future- and past-cone. Then $\phi_{i}({\rm ost}(y))$ intersects $FP(\phi_{i}(x))$. Indeed, by (1) the space-like cone from $\phi_{i}(y)$ spanned by $\phi_{i}({\rm ost}(y))$ intersects $FP(\phi_{i}(x))$, and by (2) there is an intersection point belonging to $\phi_{i}({\rm ost}(y))$. This is a contradiction with $\phi_{i}$ being future-convex.